# Learning Macroeconomic Policies based on Microfoundations: A Stackelberg Mean Field Game Approach

### v3

@@ Appendix E Additional Results
Efficiency-Equity Tradeoff of Policies
@@
In economics, the Efficiency-Equity Tradeoff is a highly debated issue. We find that our SMFG method is optimal in balancing efficiency-equity, except in cases of extreme concern for social fairness. In our study, we depict the economic efficiency (Per capita GDP) on the Y-axis and equity (wealth Gini) on the X-axis of Figure 8(a) for various policies. Different policies are represented by circles of different colors, with their sizes proportional to social welfare. Different circles of the same color correspond to different seeds. Figure 8 (a) shows that the wealth Gini indices for SMFG and AI Economist-BC are similar, but SMFG has a higher GDP, suggesting its superiority over AI Economist-BC. SMFG significantly outperforms the free market policy and AI Economist due to its higher GDPs and lower wealth Ginis. However, comparing SMFG with the Saez tax and the U.S. Federal tax policy in terms of both economic efficiency (GDP) and social equity (Gini) is challenging. Therefore, we introduce Figure 8 (b) to demonstrate the performance of different policies under various weights in a multi-objective assessment.
In Figure 8 (b), the Y-axis shows the weighted values of the multi-objective function $Y=log(\text{per capita GDP})+\alpha(\text{wealth Gini})$, and the X-axis represents the weight of the wealth Gini index. For each weight $\alpha$, we compute the multi-objective weighted values for those policies, represented as circles of different colors. Due to the logarithmic treatment of GDP in (b), when $\alpha=10$, the overall objective focuses solely on social fairness; when $\alpha=0$, the overall objective is concerned only with efficiency. Our findings in (b) reveal that only when $\alpha\geq 8$, which indicates a substantial emphasis on social equity, does the Saez tax outperform SMFG. However, SMFG consistently proves to be the most effective policy under a wide range of preference settings.
Behavior Cloning Experiments
We conduct behavior cloning based on real data to simulate the behavior strategies of households in realistic scenarios, which are then used in Experiment 5 to compare with SMFG Households. We collect the statistical data from the 2022 Survey of Consumer Finances (SCF) (https://www.federalreserve.gov/econres/scfindex.htm) as the real data buffer $\mathcal{D}_{real}$.
@@
This experiment conducts behavior cloning on networks for four different household policies: Multilayer Perceptron (MLP), AI economist’s network (MLP+LSTM+MLP), SMFG-S, and SMFG-MF network. The first two, as their network outputs, are probability distributions, use negative log-likelihood loss (Figure 9 left); the latter two’s networks employ deterministic policies, hence they use mean square error loss against real data (Figure 9 right). The loss convergence curve of behavior cloning is shown in Figure 9. It can be observed that the AI economist’s network, due to its complexity, struggles to converge to near -1 like MLP. The losses corresponding to MFRL and SMFG-MF can converge to below 0.1.
@@
Appendix F Economic Model Details
@@
where $C_{t}=\sum_{i}^{N}c^{i}_{t}$ represents the total consumption of consumers, and $G_{t}$ is government spending. The supply, demand, and price represent the states of the market.
Economic Shocks
@@ Appendix G Hyperparameters
Table 5: Hyperparameters of SMFG methods and its variants.
@@
Table 6: Hyperparameters of SMFG-MF algorithm different from SMFG method.
@@
Table 7: Hyperparameters of AI Economist Algorithm different from SMFG approach.
@@
This research introduces a novel SMFG method, designed to optimize macroeconomic policies by modeling complex interactions at the micro level. The potential impact of this work extends across several domains:
@@
In summary, the proposed SMFG framework and SMFRL algorithm hold the potential to significantly impact both academic research and practical policy making, offering a new perspective on dynamic economic modeling that prioritizes realistic, individual-level responses within large-scale economic systems.

### v4

@@ E.3 Training Curves for Various Tax Policies
E.4 Efficiency-Equity Tradeoff of Policies
@@
In economics, the Efficiency-Equity Tradeoff is a highly debated issue. We find that our DSMFG method is optimal in balancing efficiency-equity, except in cases of extreme concern for social fairness. In our study, we depict the economic efficiency (Per capita GDP) on the Y-axis and equity (wealth Gini) on the X-axis of Figure 8(a) for various policies. Different policies are represented by circles of different colors, with their sizes proportional to social welfare. Different circles of the same color correspond to different seeds. Figure 8 (a) shows that the wealth Gini indices for DSMFG and AI Economist-BC are similar, but DSMFG has a higher GDP, suggesting its superiority over AI Economist-BC. DSMFG significantly outperforms the free market policy and AI Economist due to its higher GDPs and lower wealth Ginis. However, comparing DSMFG with the Saez tax and the U.S. Federal tax policy in terms of both economic efficiency (GDP) and social equity (Gini) is challenging. Therefore, we introduce Figure 8 (b) to demonstrate the performance of different policies under various weights in a multi-objective assessment.
In Figure 8 (b), the Y-axis shows the weighted values of the multi-objective function $Y=log(\text{per capita GDP})+\alpha(\text{wealth Gini})$, and the X-axis represents the weight of the wealth Gini index. For each weight $\alpha$, we compute the multi-objective weighted values for those policies, represented as circles of different colors. Due to the logarithmic treatment of GDP in (b), when $\alpha=10$, the overall objective focuses solely on social fairness; when $\alpha=0$, the overall objective is concerned only with efficiency. Our findings in (b) reveal that only when $\alpha\geq 8$, which indicates a substantial emphasis on social equity, does the Saez tax outperform DSMFG. However, DSMFG consistently proves to be the most effective policy under a wide range of preference settings.
E.5 Behavior Cloning Experiments
We conduct behavior cloning based on real data to simulate the behavior strategies of households in realistic scenarios, which are then used in Experiment 5.4 to compare with DSMFG followers. We collect the statistical data from the 2022 Survey of Consumer Finances (SCF) (https://www.federalreserve.gov/econres/scfindex.htm) as the real data buffer $\mathcal{D}_{real}$.
@@ E.5 Behavior Cloning Experiments
This experiment conducts behavior cloning on networks for four different household policies: Multilayer Perceptron (MLP), AI economist’s network (MLP+LSTM+MLP), DSMFG w/o S, and DSMFG w/o MF network. The first two, as their network outputs, are probability distributions, use negative log-likelihood loss (Figure 9 left); the latter two’s networks employ deterministic policies, hence they use mean square error loss against real data (Figure 9 right). The loss convergence curve of behavior cloning is shown in Figure 9. It can be observed that the AI economist’s network, due to its complexity, struggles to converge to near -1 like MLP. The losses corresponding to MFRL and DSMFG w/o MF can converge to below 0.1.
@@
Appendix F TaxAI
F.1 Introduction of TaxAI
TaxAI is a novel Multi-Agent Reinforcement Learning (MARL) environment designed to model dynamic interactions among governments, households, firms, and financial intermediaries. Built on the Bewley-Aiyagari economic model, TaxAI addresses critical limitations of existing economic simulators by offering enhanced scalability, realism, and benchmarking capabilities.
  • Scalability: TaxAI simulates dynamic interactions involving up to 10,000 households, significantly surpassing the scale of prior simulators and enabling large-scale analysis.
  • Realism: Calibrated with real-world data from the 2022 Survey of Consumer Finances (SCF), TaxAI ensures its simulations reflect realistic economic scenarios, improving the relevance of its outcomes for policymaking.
  • Benchmarking: TaxAI evaluates 7 MARL algorithms against 2 traditional economic approaches (e.g., genetic algorithms, dynamic programming), demonstrating the superiority of MARL in addressing dynamic, partially observable economic environments.
  • Policy Optimization: TaxAI leverages MARL’s adaptive learning capabilities to model complex government-household interactions and discover optimal tax policies that promote growth and equity.
With its ability to integrate scalability, real-world calibration, and MARL-based adaptive optimization, TaxAI sets a new benchmark for realistic and effective economic simulators, providing actionable insights for policy design and implementation. Therefore, TaxAI is highly suitable as the experimental environment for this paper, particularly due to its scalability and realism.
F.2 Economic model details
@@
where $C_{t}=\sum_{i}^{N}c^{i}_{t}$ represents the total consumption of households, and $G_{t}$ is government spending. The supply, demand, and price represent the states of the market.
F.3 Economic Shocks
@@ Appendix G Hyperparameters
Table 6: Hyperparameters of DSMFG methods and its variants.
@@
Table 7: Hyperparameters of DSMFG w/o MF algorithm different from DSMFG method.
@@
Table 8: Hyperparameters of AI Economist Algorithm different from DSMFG approach.
@@
This research introduces a novel DSMFG method, designed to optimize macroeconomic policies by modeling complex interactions at the micro level. The potential impact of this work extends across several domains:
@@
In summary, the proposed DSMFG framework and SMFRL algorithm hold the potential to significantly impact both academic research and practical policy making, offering a new perspective on dynamic economic modeling that prioritizes realistic, individual-level responses within large-scale economic systems.